Quick Computing installed its previous generation of computer chip manufacturing equipment 3 years ago. Some of that older equipment will become unnecessary when the company goes into production of its new product. The obsolete equipment, which originally cost $36.50 million, has been depreciated straight-line over an assumed tax life of 5 years, but it can be sold now for $17.30 million. The firm’s tax rate is 30%. What is the after-tax cash flow from the sale of the equipment?
Answer #1

Annual depreciation=(cost-salvage value)/useful life

=(36.5/5)=$7.3 million/year

Hence book value as on date of sale=cost- accumulated depreciation

=36.5-(7.3*3)

=$14.6 million

Hence gain on sale=(17.3-14.6)=$2.7 million

Hence after tax cash flow=sale proceeds-(tax rate*gain on sale)

=17.3-(2.7" 0.3) million

=$16.49 million.